# Uncovering the first-infall history of the LMC through its dynamical impact in the Milky Way halo

### v1

@@ Abstract
The gravitational interactions between the LMC and the Milky Way can give rise to dynamical perturbations in the MW halo, leading to a biased distribution of stellar density and other kinematic signals. These disequilibrium phenomena exhibit variations under different parameter combinations of the MW-LMC model. In this work, we run 50 high-resolution N-body simulations spanning different masses and halo shapes of the Milky Way and LMC and investigate how the LMC-induced perturbations evolve with these model parameters. We measure the magnitude of kinematic perturbations from the mean velocities of simulated halo stars and identify a discontinuity between the first-infall and second-passage scenarios of the LMC’s orbital history. Specifically, the Galactocentric latitudinal velocity in the first-infall scenario is predicted to hovers around 13 km/s for stars at a distance of 50-100 kpc due the the dynamical influence of the LMC, while in the second-passage scenario, the expected perturbation subsides to around 6 km/s. We demonstrate that, due to the short dynamical times of the Galactic inner halo, the reduced perturbation magnitude in the second-passage scenario is mainly a result of the LMC’s second infall into the MW, which starts at a much lower velocity relative to the inner halo compared to the first-infall scenario. Using a subset of $\sim 1200$ RR Lyrae stars located in the outer halo (50 kpc $\leq R_{\mathrm{GC}}<$ 100 kpc), which are selected from a larger sample of 135,873 RR Lyrae stars with precise distance estimates from Gaia, we find the mean latitudinal velocity ($v_{b}$) in the heliocentric frame to be $\langle v_{b}\rangle=30.8\pm 4.0$ km/s. The observation contradicts the second-passage scenario and supports the first-infall scenario with a massive LMC ($\sim$ $2.1\times 10^{11}\mathrm{M}_{\odot}$) at infall, an oblate MW halo with a virial mass $M_{200}<1.4\times 10^{12}\mathrm{M}_{\odot}$ and a flattening parameter $q>0.7$. Our study opens up a new possibility to infer the orbital history and some key properties of the MW and LMC through the LMC-induced kinematic perturbations in the MW halo, paving the way for the future application of simulation-based inference in the MW-LMC system.
@@ 1 Introduction
The Large Magellanic Cloud (LMC) is the most luminous and prominent satellite of the Milky Way (MW) which has recently passed the pericenter of its orbit at about 50 kpc away from the galactic center, and has a mass only 5–10 times smaller than that of the Milky Way (e.g., Pardy et al., 2018; Laporte et al., 2018a; Erkal et al., 2019). Although interactions between the orbiting satellite galaxies and the dark matter halo of their hosts have been studied analytically since Chandrasekhar (1943), our understanding of the LMC’s orbital history and its impact on the Milky Way dynamics has only began to evolve in the last decade thanks to development of numerical simulations (e.g. Besla et al., 2010; Laporte et al., 2018a, b; Garavito-Camargo et al., 2019; Vasiliev & Belokurov, 2020) and several large-scale surveys such as Gaia (Gaia Collaboration et al., 2018), APOGEE (Majewski et al., 2017), SEGUE (Yanny et al., 2009), LAMOST (Cui et al., 2012) and H3 (Conroy et al., 2019).
@@
[…]
@@
The past orbit of the LMC might imprint different gravitational interactions between the LMC and the MW halo (Choi et al., 2009; Garavito-Camargo et al., 2021; Vasiliev, 2023a). Notably, orbiting satellite galaxies, such as the LMC, can create density asymmetries in the dark matter and stellar halos of their host galaxy, which subsequently affect the kinematic signature of the halo stars (Gómez et al., 2015; Ogiya & Burkert, 2016; Laporte et al., 2018a, b; Garavito-Camargo et al., 2019; Petersen & Peñarrubia, 2020; Erkal et al., 2021; Garavito-Camargo et al., 2021; Vasiliev, 2023b). In the MW-LMC system, the LMC induces two major responses in the MW’s halo: (a) the local effect, dynamical friction (DF) wake of the LMC, which is an overdensity of stars following the past trajectory of the LMC (Chandrasekhar, 1943; Choi et al., 2009; Ogiya & Burkert, 2016; Garavito-Camargo et al., 2019; Tamfal et al., 2021), and (b) the global effect, characterized by asymmetries in both kinematics and density of the MW on a large scale, arises mainly from the offset in net motion of the MW’s inner regions relative to its outer regions (i.e. the reflex motion of the MW) (Weinberg, 1995; Gómez et al., 2015; Garavito-Camargo et al., 2019; Erkal et al., 2021). These dynamical effects in the MW halo contain valuable information about the orbital frequency and past trajectory of the LMC.
@@ 2 Data
Figure 1: Our final 1,165 RR Lyrae stars. The on-sky spatial distribution in the Galactic longitude-latitude ($l$-$b$) space in shown in the top panel, and the distribution of the Galactic distance is shown in the botton panel. The observation sample with which we will compare with our simulations includes RR Lyrae stars within the Galactocentric radius range of 50 kpc < $R_{\mathrm{gal}}$ < 100 kpc. The sample has removed notable halo substructures like LMC, SMC, Sagittarius stream ($\left|b_{\text{Sgr}}\right|<15^{\circ}$), stars distributed near the galactic disk $\left(|b|<10^{\circ}\right)$ and 4 visible dSph clusters to better study the dynamical effect of the LMC on the Milky Way halo.
@@
Our final sample consists of 1,165 RR Lyrae stars. This is a significant reduction from the initial $15,594$ because the dominant sample of RR Lyrae stars is from the LMC and SMC, which is excluded in our case. The spatial distribution as well as the distance of our RR Lyrae sample is shown in Figure  1.
@@ 3 Simulations
Table 1: The present-day phase-space coordinates of the LMC in Galactocentric frame. All quantities are calculated from the values in Kallivayalil et al. (2013), by adopting the Solar reflex motion $V_{c,\text{ peak }}(8.29\mathrm{kpc})\approx 239$ km/s (McMillan, 2011) and solar peculiar velocity $(U,V,W)_{\odot}=\left(11.1_{-0.75}^{+0.69},12.24_{-0.47}^{+0.47},7.25_{-0.36}^{+0.37}\right)$ km/s (Schönrich et al., 2010). The standard errors on each component are calculated by Monte-Carlo sampling the errors on the observables.
@@ 3.2 Orbital reconstruction
$\ddot{\boldsymbol{x}}_{\mathrm{LMC}}=\frac{d\Phi_{\mathrm{bulge},\mathrm{MW}}}{d\boldsymbol{x}}+\frac{d\Phi_{\mathrm{disk},\mathrm{MW}}}{d\boldsymbol{x}}+\frac{d\Phi_{\mathrm{halo},\mathrm{MW}}}{d\boldsymbol{x}}+\frac{\boldsymbol{F}_{\mathrm{df}}}{\boldsymbol{M}_{\mathrm{LMC}}}$ (4)
$\ddot{\boldsymbol{x}}_{\mathrm{MW}}=\frac{d\Phi_{\mathrm{LMC}}}{d\boldsymbol{x}}$ (5)
@@ 4 Results of simulations
We now analyze the results of our simulations in this section. We first present the orbital histories of the LMC in a few representative cases which guide our selection of model parameters for further investigation. Subsequently, we conduct a comparative analysis of the LMC-induced dynamical effects in both the first-infall and second-passage scenarios. We provide an explanation for the reduced perturbation magnitudes observed in the second-passage scenario. We then quantify the magnitudes of the dynamical effects with their average values and explore how they are influenced by the key model parameters identified earlier. Finally, we compare the perturbation magnitudes in the latitudinal velocity of halo stars, located within a Galactocentric distance range of 50 to 100 kpc, between our RR Lyrae star sample and the simulated star samples derived from our full simulation grid. Our findings suggest that the magnitude of these perturbations can serve not only as an indicator of the LMC’s orbital history but also as a mean to place constraints on the properties of the halos of both the MW and the LMC. In Section 4.1, we show the past trajectories of the LMC in the N-body simulation in three representative cases. In Section 4.2, we first reproduce some of the known LMC-induced dynamical effects in the MW halo for the first-infall scenario, and then explore them in the second-passage scenario. In Section 4.3, we quantify the kinematic perturbations of halo stars with summary statistics, their mean values, and illustrate the reason for our choice. In Section 4.4, we evaluate the influence of some key parameters of the MW and the LMC on the magnitude of the kinematic perturbations in Galactocentric frame, starting from our fiducial MW-LMC model. Finally, in Section 4.5, we compare the mean latitudinal velocity in the heliocentric Galactic frame between the RR Lyrae sample and the simulated star catalogs derived from our grid of MW-LMC simulations. We then use the results of the comparison to constrain the orbital history of the LMC and some key properties of the two galaxies.
@@ 4.1 Orbital histories of the LMC
Empirical evidence suggests that the shape of the galactic halo, characterized by its flattening parameter $q$, is not necessarily spherical (e.g., Vera-Ciro & Helmi, 2013; Hernitschek et al., 2018; Hattori et al., 2021; Helmi, 2004; Bowden et al., 2016). Therefore, a natural question arises: how does the variation in $q$ affect the transition between the first-infall and second-passage scenarios of the LMC? In Figure 3, we show the LMCs’ orbital histories in an oblate MW halo, which has the same mass of galaxies ($M_{\text{MW}}=1.3\times 10^{12}\mathrm{M}_{\odot}$, $M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$) as the model in the bottom panel of Figure 2 but the MW halo is flattened in the direction perpendicular to the Galactic disk with a flattening parameter $q$ = 0.7. Due to the flattened shape of the halo, the Galactic potential at the present-day location of the LMC has increased relative to that in a spherical halo. Since the current kinematic energy of the LMC relative to the MW remains constant in both cases, this change in potential energy leads to a higher orbital energy, making the LMC less gravitationally bound to the Milky Way. Thus, compared to the case with a spherical ($q$ = 1.0) halo of equal mass, where the LMC makes two pericentric passages, the LMC’s past orbit around an oblate halo belongs to the first-infall scenario this time. This underscores the significance of treating the shape of the MW halo as a crucial variable in our simulations.
@@ 4.2 Dynamical effects in the Milky Way halo
We show the Galactocentric sky maps of density and velocity perturbations in three representative cases in Figure 4-6, using the same simulation suits described in Figure 2 ($M_{\text{MW}}=0.7,1.0,1.3\times 10^{12}\mathrm{M}_{\odot}$, $M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$, $q$ = 1.0 and velocity profile of the MW halo is isotropic $\beta(r)=0$). We show the density contrast (top row, the relative variations between the local density and the average density), the line-of-sight velocity (middle row) and the latitudinal velocity (bottom row) within the stellar halo at two radii intervals: 30-60 kpc (left column) and 60-90 kpc (right column). Maps of the inner 30 kpc of the halo are not shown in this study, as there are no significant perturbations in this region because the majority of particles within this area moving at a similar velocity to that of the Galactic disk. We note that the density and kinematic features of the first-infall scenario demonstrated in Figure 4 and Figure 5 are consistent with those obtained from first-infall simulations in previous studies (Garavito-Camargo et al., 2019; Vasiliev, 2023a). The dynamical influence of the LMC on the MW can be broadly understood as follow:
@@ 4.3 Measure the kinematic perturbations with mean values
We are also aware of other methods have been proposed in previous studies to quantitatively study the LMC-induced perturbations in density and velocity fields. For example, Cunningham et al. (2020) and Gaia Collaboration et al. (2021) investigated the spherical harmonic expansion of the velocity field and the basis function expansion of the density field within the simulated MW stellar halos derived from Garavito-Camargo et al. (2019). While these methods show the potential for disentangling the dynamical effects from the initial structure of the MW halo, the method might face more complications for our study due to the measurement uncertainties and limited sky coverage. The expansion of the density field can also be subject to the complex selection function that is beyond the scope of this study. For these reasons, we keep using simpler summary statistics and defer the discussion of more detailed distributions of density and kinematic perturbations to future studies.
@@ 4.4 Quantitative assessment of the kinematic features
Figure 9: Comparison of the mean latitudinal velocity ($v_{b}$) as calculated in the heliocentric Galactic frame between observational data and simulation results. The mean and uncertainty of the RR Lyrae stars are shown with the yellow dash-dotted line and the corresponding shaded region. For the mock data, their mean values and uncertainties are represented by the solid (first-infall), dashed (second-passage) curves and shaded regions in different colors. The blue shaded region represents MW-LMC systems with a LMC mass of $1.5\times 10^{11}\mathrm{M}_{\odot}$, featuring a spherical Galactic halo with isotropic ($\beta(r)=0$) velocity profile but varying MW halo masses. The red shaded region represents the MW-LMC models with the same MW parameters as the blue shaded region but with a heavier LMC ($M_{\text{LMC}}=2.1\times 10^{11}\mathrm{M}_{\odot}$). The green shaded region maintains the same MW and LMC potentials as the blue shaded region but assumes the MW halo has a radially anisotropic ($\beta(r)=-0.15-0.2\alpha(r)$, where $\alpha(r)=\frac{\mathrm{d}\ln\rho(r)}{d\ln r}$) velocity profile. The purple shaded region represents the cases where the shape of the MW halo in the blue shaded region transitions to an oblate shape ($q$ = 0.7), with its minor axis perpendicular to the Galactic disk. We find that the RR Lyrae sample has a mean latitudinal velocity $\langle v_{b}\rangle=30.8\pm 4.0$ km/s, which supports the first-infall scenario as the likely orbital history for the LMC. In addition, our findings favour a massive LMC model with virial mass close to $2.1\times 10^{11}\mathrm{M}_{\odot}$ at the time of infall, an oblate MW halo with the flattening parameter $q$ larger than 0.7 and a mass of the MW halo less than $1.4\times 10^{12}\mathrm{M}_{\odot}$ (for detailed discussion, see Section 5.6).
Varying mass of the MW halo: On the other hand, models with a constant LMC mass but varying MW halo masses (middle column) demonstrate a shift in the LMC’s orbital history from the first-infall scenario to the second-passage scenario when the MW mass exceeds $1.0\times 10^{12}\mathrm{M}_{\odot}$. Following this transition, the magnitude of kinematic signals decreases by about a factor of two, resulting in a discontinuity between the first-infall and second-passage scenarios. The underlying cause of this transition is the reduced gravitational potential at the LMC’s present-day location in a more massive halo. Given that the LMC’s current kinematic energy relative to the MW remains unchanged regardless of variations in halo mass, this decrease in potential energy results in a lower orbital energy and the LMC becomes more gravitationally bound to the Milky Way.
@@
Varying flattening parameter of the MW halo: Moreover, the orbital history of the LMC is also sensitive to the variation in the shape of the MW halo. Models with the same MW and LMC mass ($M_{\text{MW}}=1.0\times 10^{12}\mathrm{M}_{\odot}$, $M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$) but different flattening parameters $q$ (ranging from 0.5 to 1.3, as shown in the right column) show a discontinuity in the magnitude of kinematic perturbations when transitioning from $q$leq 1.0 to $q$ = 1.1. This transition suggests that, for this MW mass, the orbital history of the LMC aligns with the first-infall scenario when the MW halo is oblate or spherical. However, it can shift to the second-passage scenario when the MW halo becomes prolate. We present a discussion on the degeneracy between halo mass and shape when it comes to determining the orbital history of the LMC in section 5.6. Similar to the transition associated with variations in the MW halo mass, the gravitational potential at the LMC’s present-day location is higher in an oblate halo compared to a prolate halo. Given that the LMC’s current kinematic energy relative to the MW remains unchanged in both cases, this variation in potential energy results in the LMC having a higher orbital energy and being less gravitationally bound when orbiting an oblate halo, whereas it possesses a lower orbital energy and is more gravitationally bound when orbiting a prolate halo.

### v2

@@ Abstract
The gravitational interactions between the LMC and the Milky Way cause dynamical perturbations in the MW halo, leading to biased distributions of stellar density and kinematics. We run 50 high-resolution N-body simulations exploring varying masses and halo shapes of the MW and LMC to study the evolution of LMC-induced perturbations. By measuring mean velocities of simulated halo stars, we identify a discontinuity between the first-infall and second-passage scenarios of the LMC’s orbital history. In the first-infall, the Galactocentric latitudinal velocity hovers around 16 km/s for stars at 50-100 kpc, while it subsides to about 8 km/s in the second-passage scenario. We demonstrate that, this reduced perturbation magnitude in the second-passage scenario is mainly due to the short dynamical times of the Galactic inner halo and the lower velocity of the LMC during its second infall into the MW. Using a subset of $\sim 1100$ RR Lyrae stars located in the outer halo (50 kpc $\leq R_{\mathrm{GC}}<$ 100 kpc) with precise distance estimates from Gaia, we find the mean latitudinal velocity ($v_{b}$) in the Galactocentric frame to be $\langle v_{b}\rangle=18.1\pm 4.1$ km/s. The observation supports the first-infall scenario with a massive LMC ($\sim$ $2.1\times 10^{11}\mathrm{M}_{\odot}$) at infall, an oblate MW halo with a virial mass $M_{200}<1.4\times 10^{12}\mathrm{M}_{\odot}$ and a flattening parameter $q>0.7$. Our study indicates that LMC-induced kinematic disturbances can reveal its orbital history and key characteristics, as well as those of the MW. This approach shows promise in helping determine fundamental parameters of both galaxies.
@@ 1 Introduction
The Large Magellanic Cloud (LMC) is the most luminous and prominent satellite of the Milky Way (MW) which has recently passed the pericenter of its orbit at about 50 kpc away from the galactic center, and has a mass only 5–10 times smaller than that of the Milky Way (e.g., Kallivayalil et al., 2013; Pardy et al., 2018; Laporte et al., 2018a; Erkal et al., 2019). Although interactions between the orbiting satellite galaxies and the dark matter halo of their hosts have been studied analytically since Chandrasekhar (1943), our understanding of the LMC’s orbital history and its impact on the Milky Way dynamics has only began to evolve in the last decade thanks to development of numerical simulations (e.g. Besla et al., 2010; Gómez et al., 2015; Laporte et al., 2018a; Laporte et al., 2018b; Garavito-Camargo et al., 2019; Vasiliev et al., 2021) and several large-scale surveys such as Gaia (Gaia Collaboration et al., 2018), APOGEE (Majewski et al., 2017), SEGUE (Yanny et al., 2009), LAMOST (Cui et al., 2012) and H3 (Conroy et al., 2019).
@@
  • The LMC’s star formation rate has been unusually low until a recent burst starting about 3–4 Gyr ago (e.g., Harris & Zaritsky, 2009; Meschin et al., 2014; Massana et al., 2022), which could have been triggered by the compression of gas when the LMC entered the MW, supporting the first-infall scenario of the LMC’s orbital history. However, this phenomenon can also be attributed to interactions with the Small Magellanic Cloud (SMC), as a synchronised ignition of star formation is found in the star formation histories of both the LMC and SMC around the same time period (Massana et al., 2022).
@@
The past orbit of the LMC might imprint different gravitational interactions between the LMC and the MW halo (Choi et al., 2009; Garavito-Camargo et al., 2021; Vasiliev, 2023a). Notably, orbiting satellite galaxies, such as the LMC, can create density asymmetries in the dark matter and stellar halos of their host galaxy, which subsequently affect the kinematic signature of the halo stars (Gómez et al., 2015; Ogiya & Burkert, 2016; Laporte et al., 2018a; Laporte et al., 2018b; Garavito-Camargo et al., 2019; Petersen & Peñarrubia, 2020; Erkal et al., 2021; Garavito-Camargo et al., 2021; Vasiliev, 2023b). In the MW-LMC system, the LMC induces two major responses in the MW’s halo: (a) the local effect, dynamical friction (DF) wake of the LMC, which is an overdensity of stars following the past trajectory of the LMC (Chandrasekhar, 1943; Choi et al., 2009; Ogiya & Burkert, 2016; Garavito-Camargo et al., 2019; Tamfal et al., 2021), and (b) the global effect, characterized by asymmetries in both kinematics and density of the MW on a large scale, arises mainly from the offset in net motion of the MW’s inner regions relative to its outer regions (i.e. the reflex motion of the MW) (Weinberg, 1995; Gómez et al., 2015; Garavito-Camargo et al., 2019; Erkal et al., 2021). These dynamical effects in the MW halo contain valuable information about the orbital frequency and past trajectory of the LMC.
@@ 2 Data
Figure 1: Our final 1,126 RR Lyrae stars. The on-sky spatial distribution in the Galactic longitude-latitude ($l$-$b$) space in shown in the top panel, and the distribution of the Galactic distance is shown in the botton panel. The observation sample with which we will compare with our simulations includes RR Lyrae stars within the Galactocentric radius range of 50 kpc < $R_{\mathrm{gal}}$ < 100 kpc. The sample has removed notable halo substructures like LMC, SMC, Sagittarius stream ($\left|b_{\text{Sgr}}\right|<15^{\circ}$), stars distributed near the galactic disk $\left(|b|<10^{\circ}\right)$ and 4 visible dSph clusters to better study the dynamical effect of the LMC on the Milky Way halo.
@@
Our final sample consists of 1,126 RR Lyrae stars. This is a significant reduction from the initial $15,594$ because the dominant sample of RR Lyrae stars is from the LMC and SMC, which is excluded in our case. The spatial distribution as well as the distance of our RR Lyrae sample is shown in Figure  1.
@@ 3 Simulations
Table 1: The present-day phase-space coordinates of the LMC in Galactocentric cartesian frame. All quantities are calculated from the measured values in Kallivayalil et al. (2013), by adopting the Solar reflex motion $V_{c,\text{ peak }}(8.29\mathrm{kpc})\approx 239$ km/s (McMillan, 2011) and solar peculiar velocity $(U,V,W)_{\odot}=\left(11.1_{-0.75}^{+0.69},12.24_{-0.47}^{+0.47},7.25_{-0.36}^{+0.37}\right)$ km/s (Schönrich et al., 2010). The standard errors on each component are calculated by Monte-Carlo sampling the errors on the measured proper motions [^2] Since the proper motion measurements of the LMC do not include covariances, here we only compare the reconstructed coordinates from N-body simulations with the marginal distributions of individual phase-space measurements in the Galactocentric cartesian frame., radial velocities, and distances.
@@ 3.2 Orbital reconstruction
$\ddot{\boldsymbol{x}}_{\mathrm{LMC}}=\frac{-d\Phi_{\mathrm{bulge},\mathrm{MW}}}{d\boldsymbol{x}}+\frac{-d\Phi_{\mathrm{disk},\mathrm{MW}}}{d\boldsymbol{x}}+\frac{-d\Phi_{\mathrm{halo},\mathrm{MW}}}{d\boldsymbol{x}}+\frac{\boldsymbol{F}_{\mathrm{df}}}{\boldsymbol{M}_{\mathrm{LMC}}}$ (4)
$\ddot{\boldsymbol{x}}_{\mathrm{MW}}=\frac{-d\Phi_{\mathrm{LMC}}}{d\boldsymbol{x}}$ (5)
@@ 4 Results of simulations
We now analyze the results of our simulations in this section. We first present the orbital histories of the LMC in a few representative cases which guide our selection of model parameters for further investigation. Subsequently, we conduct a comparative analysis of the LMC-induced dynamical effects in both the first-infall and second-passage scenarios. We provide an explanation for the reduced perturbation magnitudes observed in the second-passage scenario. We then quantify the magnitudes of the dynamical effects with their average values and explore how they are influenced by the key model parameters identified earlier. Finally, we compare the perturbation magnitudes in the latitudinal velocity of halo stars, located within a Galactocentric distance range of 50 to 100 kpc, between our RR Lyrae star sample and the simulated star samples derived from our full simulation grid. Our findings suggest that the magnitude of these perturbations can serve not only as an indicator of the LMC’s orbital history but also as a mean to place constraints on the properties of the halos of both the MW and the LMC. In Section 4.1, we show the past trajectories of the LMC in the N-body simulation in three representative cases. In Section 4.2, we first reproduce some of the known LMC-induced dynamical effects in the MW halo for the first-infall scenario, and then explore them in the second-passage scenario. In Section 4.3, we quantify the kinematic perturbations of halo stars with summary statistics, their mean values, and illustrate the reason for our choice. In Section 4.4, we evaluate the influence of some key parameters of the MW and the LMC on the magnitude of the kinematic perturbations in Galactocentric frame, starting from our fiducial MW-LMC model. Finally, in Section 4.5, we compare the mean latitudinal velocity in the Galactocentric frame between the RR Lyrae sample and the simulated star catalogs derived from our grid of MW-LMC simulations. We then use the results of the comparison to constrain the orbital history of the LMC and some key properties of the two galaxies.
@@ 4.1 Orbital histories of the LMC
Empirical evidence suggests that the shape of the galactic halo, characterized by its flattening parameter $q$, is not necessarily spherical (e.g., Vera-Ciro & Helmi, 2013; Hernitschek et al., 2018; Hattori et al., 2021; Helmi, 2004; Bowden et al., 2016). Therefore, a natural question arises: how does the variation in $q$ affect the transition between the first-infall and second-passage scenarios of the LMC? In Figure 3, we show the LMCs’ orbital histories in an oblate MW halo, which has the same mass of galaxies ($M_{\text{MW}}=1.3\times 10^{12}\mathrm{M}_{\odot}$, $M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$) as the model in the bottom panel of Figure 2 but the MW halo is flattened in the direction perpendicular to the Galactic disk with a flattening parameter $q$ = 0.7. Due to the flattened shape of the halo, the mass of the Milky Way enclosed within a given radius (e.g., the inner 100 kpc) decreases by about 30%. As a result, the Galactic potential at the present-day location of the LMC has increased relative to that in a spherical halo. Since the current kinematic energy of the LMC relative to the MW remains constant in both cases, this change in potential energy leads to a higher orbital energy, making the LMC less gravitationally bound to the Milky Way. Thus, compared to the case with a spherical ($q$ = 1.0) halo of equal mass, where the LMC makes two pericentric passages, the LMC’s past orbit around an oblate halo belongs to the first-infall scenario this time. This underscores the significance of treating the shape of the MW halo as a crucial variable in our simulations.
@@ 4.2 Dynamical effects in the Milky Way halo
We show the Galactocentric sky maps of density and velocity perturbations in three representative cases in Figure 4-6, using the same simulation suits described in Figure 2 ($M_{\text{MW}}=0.7,1.0,1.3\times 10^{12}\mathrm{M}_{\odot}$, $M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$, $q$ = 1.0 and velocity profile of the MW halo is isotropic $\beta(r)=0$). We show the density contrast (top row, the relative variations between the local density and the average density), the line-of-sight velocity (middle row) and the latitudinal velocity (bottom row) within the stellar halo at two radii intervals: 30-60 kpc (left column) and 60-90 kpc (right column). Maps of the inner 30 kpc of the halo are not shown in this study, as there are no significant perturbations in this region because the majority of particles within this area moving at a similar velocity to that of the Galactic disk. We note that the density and kinematic features of the first-infall scenario demonstrated in Figure 4 and Figure 5 are consistent with those obtained from first-infall simulations in previous studies (Garavito-Camargo et al., 2019; Erkal et al., 2020; Vasiliev, 2023a). The dynamical influence of the LMC on the MW can be broadly understood as follow:
@@ 4.3 Measure the kinematic perturbations with mean values
We are also aware of other methods have been proposed in previous studies to quantitatively study the LMC-induced perturbations in density and velocity fields. For example, Cunningham et al. (2020) and Garavito-Camargo et al. (2021) investigated the spherical harmonic expansion of the velocity field and the basis function expansion of the density field within the simulated MW stellar halos derived from Garavito-Camargo et al. (2019). While these methods show the potential for disentangling the dynamical effects from the initial structure of the MW halo, the method might face more complications for our study due to the measurement uncertainties and limited sky coverage. The expansion of the density field can also be subject to the complex selection function that is beyond the scope of this study. For these reasons, we keep using simpler summary statistics and defer the discussion of more detailed distributions of density and kinematic perturbations to future studies.
@@ 4.4 Quantitative assessment of the kinematic features
Figure 9: Comparison of the mean latitudinal velocity ($v_{b}$) as calculated in the Galactocentric frame between observational data and simulation results. The mean and uncertainty of the RR Lyrae stars are shown with the yellow dash-dotted line and the corresponding shaded region. For the mock data, their mean values and uncertainties are represented by the solid (first-infall), dashed (second-passage) curves and shaded regions in different colors. The blue shaded region represents MW-LMC systems with a LMC mass of $1.5\times 10^{11}\mathrm{M}_{\odot}$, featuring a spherical Galactic halo with isotropic ($\beta(r)=0$) velocity profile but varying MW halo masses. The red shaded region represents the MW-LMC models with the same MW parameters as the blue shaded region but with a heavier LMC ($M_{\text{LMC}}=2.1\times 10^{11}\mathrm{M}_{\odot}$). The green shaded region maintains the same MW and LMC potentials as the blue shaded region but assumes the MW halo has a radially anisotropic ($\beta(r)=-0.15-0.2\alpha(r)$, where $\alpha(r)=\frac{\mathrm{d}\ln\rho(r)}{d\ln r}$) velocity profile. The purple shaded region represents the cases where the shape of the MW halo in the blue shaded region transitions to an oblate shape ($q$ = 0.7), with its minor axis perpendicular to the Galactic disk. We find that the RR Lyrae sample has a mean latitudinal velocity $\langle v_{b}\rangle=18.1\pm 4.1$ km/s, which supports the first-infall scenario as the likely orbital history for the LMC. In addition, our findings favour a massive LMC model with virial mass close to $2.1\times 10^{11}\mathrm{M}_{\odot}$ at the time of infall, an oblate MW halo with the flattening parameter $q$ larger than 0.7 and a mass of the MW halo less than $1.4\times 10^{12}\mathrm{M}_{\odot}$ (for detailed discussion, see Section 5.6).
Varying mass of the MW halo: On the other hand, models with a constant LMC mass but varying MW halo masses (middle column) demonstrate a shift in the LMC’s orbital history from the first-infall scenario to the second-passage scenario when the MW mass exceeds $1.0\times 10^{12}\mathrm{M}_{\odot}$. Following this transition, the magnitude of kinematic signals decreases by about a factor of two, resulting in a discontinuity between the first-infall and second-passage scenarios. The underlying cause of this transition is the increased mass enclosed within a given radius from the Galactic center, which leads to a reduced gravitational potential at the present-day location of the LMC in a more massive halo. Given that the LMC’s current kinematic energy relative to the MW remains unchanged regardless of variations in halo mass, this decrease in potential energy results in a lower orbital energy and the LMC becomes more gravitationally bound to the Milky Way.
@@
Varying flattening parameter of the MW halo: Moreover, the orbital history of the LMC is also sensitive to the variation in the shape of the MW halo. Models with the same MW and LMC mass ($M_{\text{MW}}=1.0\times 10^{12}\mathrm{M}_{\odot}$, $M_{\text{LMC}}=1.5\times 10^{11}\mathrm{M}_{\odot}$) but different flattening parameters $q$ (ranging from 0.5 to 1.3, as shown in the right column) show a discontinuity in the magnitude of kinematic perturbations when transitioning from $q$leq 1.0 to $q$ = 1.1. This transition suggests that, for this MW mass, the orbital history of the LMC aligns with the first-infall scenario when the MW halo is oblate or spherical. However, it can shift to the second-passage scenario when the MW halo becomes prolate. We present a discussion on the degeneracy between halo mass and shape when it comes to determining the orbital history of the LMC in section 5.6. Similar to the transition associated with variations in the MW halo mass, the enclosed mass within a given radius is lower in an oblate halo compared to a prolate halo. In our cases, the enclosed mass within the inner 100 kpc increases by 22% when the flattening parameter changes from 0.9 to 1.1. Consequently, the gravitational potential at the LMC’s present-day location is significantly higher in an oblate halo. Given that the LMC’s current kinematic energy relative to the MW remains unchanged in both cases, this variation in potential energy results in the LMC having a higher orbital energy and being less gravitationally bound when orbiting an oblate halo, whereas it possesses a lower orbital energy and is more gravitationally bound when orbiting a prolate halo.